OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

Are the Equations X-y=-3 and x+3y=5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

simply subtract both oof them you will get...-4y=-8=>y=2 put it in eq 1 you will get x-2=-3 => x=-3+1 => x=-1 so sol set is (x,y)=(-1,2)

OpenStudy (anonymous):

so x=-1

OpenStudy (anonymous):

thanks, can you help with a few more? they're about square roots and stuff and i've never really understood those

OpenStudy (anonymous):

Okay, so For the First equation, solve for X. and you will have x=y-3. Plug that equation into the second equation. and you will get x=-1/2.

OpenStudy (anonymous):

Oh. Or Not.

OpenStudy (anonymous):

What is 2 to the power of 4 over 3 equal to? ;-;

OpenStudy (anonymous):

Oh, Well. That can be turned into a square root.\[\sqrt[3]{2}^4\]

OpenStudy (anonymous):

\[3\sqrt8, 2\sqrt8, 3\sqrt16, and 2\sqrt16\]

OpenStudy (anonymous):

those are the choices they gave me .-.

OpenStudy (anonymous):

its 3sqrt16 simplt put 2^4=16 in the sqrt...

OpenStudy (anonymous):

y=2.x=-1

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

medal plz lol

OpenStudy (anonymous):

can i get help on one more?

OpenStudy (anonymous):

y not..but ask in new question box..after closing it and mention me there...

OpenStudy (anonymous):

@kittycatgoesmoo okay..?

OpenStudy (anonymous):

okay sorry, it's my first time using this

OpenStudy (anonymous):

its okay...